# The GECKOS Survey: Identifying kinematic sub-structures in edge-on galaxies

### v1

@@ 4.2 Evidence for nuclear discs
From Figure 16a, it appears that the peak $h_{3}$ value is correlated with the physical size of the $h_{3}$ vs. radius ‘zig-zag’ feature for the four galaxies hosting nuclear disc signatures. We observe a sharp transition between a negative $h_{3}$ gradient with radius and a positive gradient in the central regions (corresponding to a sharp transition between $h_{3}$-$V_{\star}$ anti-correlation and correlation), which Chung & Bureau (2004) attributed to a dense and (quasi-)axisymmetric central stellar disc. We therefore measure the size of the nuclear discs detected in the four GECKOS galaxies as this transition point. We note that we could also have measured the sizes from the transitions in $h_{4}$, or the bumps in the rotation curves, but found that the signal was clearest in $h_{3}$. Gadotti et al. (2020) does similar but using the $V_{\star}/\sigma_{\star}$ peak, terming this size the kinematic radius, or $r_{\rm{k}}$. We note here that these techniques of measuring nuclear disc size will not return the full extent of the nuclear disc, just the (light-weighted) region where they dominate over background stellar populations. In all likelihood, they will be larger than these kinematically-defined sizes. We list the nuclear disc sizes and peak $h_{3}$ values for the four GECKOS galaxies that exhibit signs of nuclear disc presence in Table 2, noting that the sizes obtained are in line with those measured for the TIMER survey by Gadotti et al. (2020) and de Sá-Freitas et al. (2023).
PGC 044931 possesses the strongest peak in $h_{3}$, and the largest $r_{\rm{k}}$ value, which we expect to be the brightest (and possibly largest) nuclear disc. Interestingly, while NGC 3957 displays the second greatest $h_{3}$ peak, it is IC 1711 with the larger $r_{\rm{k}}$. If we interpret greater negative and positive $h_{3}$ to be excess material at approaching and receding velocities in the LOSVD (beyond that of a pure Gaussian), respectively, then we can assume that there are either more bright stars (or many more faint stars) in the nuclear disc of NGC 3957, or the difference in rotation speed is greater between the nuclear disc and main disc of NGC 3957 than IC 1711. Alternatively, we might be observing geometrical differences in nuclear disc orientations. Given that $x_{2}$ orbits are often somewhat elongated, and aligned perpendicular to the direction of elongation of the bar (Contopoulos & Papayannopoulos, 1980), the measured projected radius of the nuclear disc may vary as a function of bar viewing angle. In this manner, the combination of peak $h_{3}$ and $r_{\rm{k}}$ may be able to be used to infer bar viewing angles. Simulations should be employed to further investigate this idea.
@@ 4.3.1 Classical bulges
A common predictor for the presence of a visually compact spheroidal structure at the centre of a galaxy (often referred to as a classical bulge, and expected to be dynamically hotter than surrounding components e.g. Kormendy & Kennicutt, 2004) is a high central value for $\sigma_{\star}$. Méndez-Abreu et al. (2014), for example, defined any galaxy with $\sigma_{\star}>200\leavevmode\nobreak\ \rm{km}\leavevmode\nobreak\ {s}^{-1}$ to possess a classical bulge. Based on an extensive literature review, Fisher & Drory (2016) updated this value to $\sigma_{\star}>130\leavevmode\nobreak\ \rm{km}\leavevmode\nobreak\ {s}^{-1}$. Other works combine photometric and spectroscopic indicators, for example Neumann et al. (2017) measure the inner slope of the radial velocity dispersion profile within the photometrically-defined bulge region. Several GECKOS galaxies possess a peak in their dispersion profile in their central regions, and for IC 1711, this peak value is $\sim 130$ km s-1. Following the literature, we would then attribute this peak to a classical bulge, however, there are also alternate explanations.
@@
Returning to the case of IC 1711, and also IC 5244 (with peak $\sigma_{\star}\sim 110$ km s-1), despite their high $\sigma_{\star}$ peak, we conclude that the BP bulge morphology, clear kinematic nuclear disc signatures, and additional high-order kinematic information point to likely end-on bars being the cause of the $\sigma_{\star}$ behaviour in these galaxies. The observations can be fully explained by the presence of an end-on bar, without the need to invoke a classical bulge. For this reason, we can say that we see no signs of classical bulges in the 12 galaxies in this sample. It will be informative to examine the full GECKOS sample for such features, though we note that the Milky Way itself possesses cylindrical rotation (Shen et al., 2010) and structural and kinematic properties (Bland-Hawthorn & Gerhard, 2017) inconsistent with the presence of a classical bulge.
@@ 4.3.2 Kinematic disc structures
With this, we have explained the kinematic behaviour of 11/12 GECKOS galaxies (excluding UGC 00903, whose kinematics are considerably more complex and require detailed modelling) by a combination of bars, BP bulges, nuclear discs, and simple disc structure(s). We see no physically-motivated reason to invoke classical bulges in this sample, a conclusion also drawn by Bittner et al. (2020) for the TIMER sample. While certainly not statistical nor representative, the GECKOS sample does span a range of bulge morphologies and bulge-to-total ratios derived from two-component decompositions. The full sample in combination with dynamical modelling will provide excellent insight into the frequency but also necessity (or lack thereof) of compact spheroidal bulges ($\sim$100s of pc) in the local Universe.
@@
ESO 484-036 Yes No Yes Slightly elongated along midplane Intermediate
ESO 079-003 Yes No Yes Slightly elongated along midplane Intermediate
ESO 120-016 Maybe No Maybe Slightly elongated along midplane Intermediate (if barred)
@@
NGC 5775 No No Slightly elongated along midplane Unbarred
5 Summary & Conclusions
To understand the diversity of kinematic sub-structures in Milky Way-mass galaxies, we examine the stellar kinematics of the first 12 targets observed as part of the GECKOS Survey of nearby edge-on galaxies. After first visually examining optical and mid-IR images to identify BP bulge structures, we examine both the 1D radial profiles and 2D maps of line-of-sight mean velocity ($V_{\star}$), velocity dispersion ($\sigma_{\star}$), and high-order skew ($h_{3}$) and kurtosis ($h_{4}$). Eight of the 12 galaxies possess boxy and/or peanut-shaped central structures, present in both $r$-band optical and 3.6-$\mu$m Spitzer imaging.
@@ 5 Summary & Conclusions
  • All galaxies show $h_{3}-V_{\star}$ anti-correlations in outer disc regions consistent with a (quasi-)axisymmetric, rotating disc of stars.
  • All galaxies that show evidence for a clear boxy-peanut shaped structure in images also possess kinematic indicators of a bar. In particular, all of these galaxies show extended regions of $h_{3}$ correlation with $V_{\star}$ permeating up to 2 kpc off the disc midplane. These features are not present in the four non-BP bulge galaxies.
  • Four BP bulge galaxies also host strong kinematic signatures of nuclear discs, including a ‘double hump’ rotation curve, central $\sigma_{\star}$ minima, $h_{3}-V_{\star}$ anti-correlation in the central regions, and strong $h_{4}$ maxima co-spatial with the expected location of the nuclear disc itself
Comparing to the 2D maps produced by N-body simulations of Iannuzzi & Athanassoula (2015), there is an array of central velocity dispersions, and varying degrees of $h_{3}$-$V_{\star}$correlation across the bar-dominated regions. In IC 5244, IC 1711, NGC 3957, and PGC 044931, there is strong evidence for the existence of nuclear discs, in the form of central $\sigma_{\star}$ depressions (NGC 3957 and PGC 044931), and central $h_{3}$-$V_{\star}$ anti-correlations (in all four galaxies). Dust obscuration and resolution effects can significantly hinder the recovery of the stellar kinematics in the central regions of edge-on galaxies, so we cannot rule out the possibility that a nuclear disc is present in all of the BP bulge galaxies; we may simply not be able to detect them.
@@
6 Author contribution statement
This project was devised by AFM, with analysis completed by AFM, and some scripts provided by JvdS. Comments were provided by the whole author list. GECKOS data reduction is led by JvdS and AFM, with assistance from EE, LC, BMC, TM, and BC. nGIST development is led by AFM and JvdS, with contributions and testing from TB, AW, ZW, LSL, LC, FP, MM, DG, EE, and CdSF.
@@ Acknowledgements.
Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 110.24AS. We wish to thank the ESO staff, and in particular the staff at Paranal Observatory, for carrying out the GECKOS observations. The authors wish to thank Adrian Bittner for useful conversations, and his blessing to continue the development of the GIST  pipeline. We also thank Lodovico Coccato for useful conversations relating to MUSE data reduction and analysis, Alex Vazdekis for his permission to distribute the MILES models with the nGIST package, and Michele Cappellari for useful conversations related to pPXF usage. This research has made use of the NASA/IPAC Extragalactic Database (NED; https://ned.ipac.caltech.edu/) operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. Part of this research was conducted by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013. DAG is supported by STFC grant ST/X001075/1. This research used the Canadian Advanced Network For Astronomy Research (CANFAR) operated in partnership by the Canadian Astronomy Data Centre and The Digital Research Alliance of Canada with support from the National Research Council of Canada the Canadian Space Agency, CANARIE and the Canadian Foundation for Innovation. TAD acknowledges support from the UK Science and Technology Facilities Council through grants ST/S00033X/1 and ST/W000830/1. FP acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (MCIN/AEI/ 10.13039/501100011033) under grant (PID2021-128131NB-I00) and the European Regional Development Fund (ERDF) ”A way of making Europe”. FP acknowledges support also from the Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant “TraNSLate” No 101108180. FF is supported by a UKRI Future Leaders Fellowship (grant no. MR/X033740/1). AP acknowledges support from the Hintze Family Charitable Foundation. LSL acknowledges CAPES for the support through grant 88887.637633/2021-00 and CNPQ through grant 200469/2022-3 LMV acknowledges support by the German Academic Scholarship Foundation (Studienstiftung des deutschen Volkes) and the Marianne-Plehn-Program of the Elite Network of Bavaria. LC and ABW acknowledge support from the Australian Research Council Discovery Project funding scheme (DP210100337) J.F-B acknowledges support from the PID2022-140869NB-I00 grant from the Spanish Ministry of Science and Innovation. MM acknowledges support from the UK Science and Technology Facilities Council through grant ST/Y002490/1.
@@ Appendix A The nGIST pipeline
The need for further development arose from both the increased computational and scientific requirements of recent IFS surveys. nGIST pushes beyond current limits in terms of memory requirements and deals better with longer optical wavelength ranges and sky residuals that are particularly problematic at redder wavelengths ($\lambda>7000\AA$). Building on the foundations of the GIST pipeline, nGIST also incorporates elements of pipelines developed for recent IFS surveys, including the SAMI stellar kinematics pipeline (van de Sande et al. 2017), and the PHANGS data analysis pipeline (DAP; Emsellem et al. 2022), which in turn is rooted in portions of the MaNGA DAP (Westfall et al. 2019). nGIST represents a significant upgrade in both scientific output and computational performance over GIST.
@@
nGIST commenced from the final release of the GIST  pipeline, v3.1.0777We note that the final release of GIST  requires python $<$ v3.6, which is increasingly deprecated and difficult to install on modern machines via package managers such as conda. At the time of writing, nGIST has been tested up to python v3.11., and has continued to be developed by adding new functionality, and updates that are appropriate for increasingly complex IFS datasets. nGIST is publicly available via a GitHub repository, https://github.com/geckos-survey/ngist, with documentation available at https://geckos-survey.github.io/gist-documentation/. It has been thoroughly tested for MUSE data, but can be employed with any galaxy IFS dataset. Below, we detail the nGIST philosophy, workflow, and major changes introduced since the last version of GIST. For a full description of the original GIST pipeline, we refer the reader to Bittner et al. (2019).
@@
$bias=0.7\times\sqrt{500N_{\lambda pixels}},$ (3)
@@
In Figure 18, we show the comparison of the stellar kinematic maps from GIST and nGIST for GECKOS iDR1 galaxy IC1711, with a target $S/N=100$ and minimum spaxel $S/N=5$. For this test, we adopted the v11.0 EMILES Base models with the BaSTI scaled-solar isochrones, that can be used in both GIST and nGIST. For these fits, the wavelength range was set to 4800–8900Å, we assumed a constant noise spectrum, and fixed the pPXF bias value to 0.14.
There is a good agreement in Figure 18 for the results from GIST (left column) and nGIST (middle column) for all kinematic parameters, though there are noticeable differences in the outer bins where the contribution from the sky lines and telluric absorption features become more dominant. In particular $h_{3}$ and $h_{4}$ show reach more extreme values in GIST. A one-to-one comparison of the results are presented in Figure 19, where the data are colour-coded by the S/N calculated from the residual of the data minus the best-fit.
@@
nGIST includes an additional spectral cleaning step, and the regions masked as a result of this step are shown as shaded pink regions. While the sky line masks are extensive (based on the UVES sky emission atlas, where every sky line of flux $>5\times 10^{-16}$ erg s-1 Å-1 cm-2 arcsec-2 was masked), some additional masks are always applied by the nGIST cleaning step. The number of masked spectral pixels as a result of the cleaning step increases as spectrum S/N decreases, and there are clear regions in the outer bin spectrum that benefit from the additional masking, particularly a regionaround $\lambda\sim 8700\AA$, which the Calcium triplet sits on top of at the redshift of this galaxy.
@@
We compare the mean stellar population age and metallicity from GIST and nGIST in Figures 23 and 24. Similar to Pessa et al. (2023), for this comparison we restrict the wavelength region to 4800-7000Å, and adopt a multiplicative polynomial of order 12. We apply minimal regularisation to the resultant mean stellar age and metallicity maps. The new cleaning algorithm in nGIST has a significant impact on the derived age and metallicity, most noticeable at large radii and scale heights, where the sky line residuals and telluric correction are most pronounced. A full test of the impact of clipping outliers, and the dependence of wavelength range on the mean stellar population parameters, is beyond the scope of this Paper, but will be explored in future work.

### v2

@@ 4.2 Evidence for nuclear discs
From Figure 16a, it appears that the magnitude of the peak $h_{3}$ value is correlated with the physical distance between the $h_{3}$ minima and maxima for the four galaxies hosting nuclear disc signatures. The $h_{3}$ extrema correspond to changes in the sign of the gradient of $h_{3}$ with radius, which also corresponds to the turnover in peak $V_{\star}/\sigma$ values used to measure nuclear disc size for the TIMER galaxy sample, and termed $r_{\rm{k}}$ (Gadotti et al., 2020). We therefore measure the size of the nuclear discs detected in the four GECKOS galaxies as half of the physical distance between $h_{3}$ extrema, terming this value $r_{h3}$. We note that we could also have measured the sizes from the transitions in $h_{4}$, or the bumps in the rotation curves, but found that the signal was clearest in $h_{3}$. We note here that these techniques of measuring nuclear disc size will not return the full extent of the nuclear disc, just the (light-weighted) region where they dominate over background stellar populations. In all likelihood, they will be larger than these kinematically-defined sizes. There are also other techniques for measuring the size of the disc, for example measuring the radius for which $h_{3}$ reaches zero again after its peak (which Chung & Bureau (2004) attributed to a dense and (quasi-)axisymmetric central stellar disc). However, to maximise comparison with recent observational works such as Gadotti et al. (2020), we choose to stick to the above definition of $r_{h3}$. We list the nuclear disc sizes and peak $h_{3}$ values for the four GECKOS galaxies that exhibit signs of nuclear disc presence in Table 2, noting that the sizes obtained are in line with those measured for the TIMER survey by Gadotti et al. (2020) and de Sá-Freitas et al. (2023).
PGC 044931 possesses the strongest peak in $h_{3}$, and the largest $r_{h3}$ value, which we expect to be the brightest (and possibly largest) nuclear disc. Interestingly, while NGC 3957 displays the second greatest $h_{3}$ peak, it is IC 1711 with the larger $r_{h3}$. If we interpret greater negative and positive $h_{3}$ to be excess material at approaching and receding velocities in the LOSVD (beyond that of a pure Gaussian), respectively, then we can assume that there are either more bright stars (or many more faint stars) in the nuclear disc of NGC 3957, or the difference in rotation speed is greater between the nuclear disc and main disc of NGC 3957 than IC 1711. Alternatively, we might be observing geometrical differences in nuclear disc orientations. Given that $x_{2}$ orbits are often somewhat elongated, and aligned perpendicular to the direction of elongation of the bar (Contopoulos & Papayannopoulos, 1980), the measured projected radius of the nuclear disc may vary as a function of bar viewing angle. In this manner, the combination of peak $h_{3}$ and $r_{h3}$ may be able to be used to infer bar viewing angles. Simulations should be employed to further investigate this idea.
@@ 4.3.1 Classical bulges
A common predictor for the presence of a visually compact spheroidal structure at the centre of a galaxy (often referred to as a classical bulge, and expected to be dynamically hotter than surrounding components e.g. Kormendy & Kennicutt, 2004) is a high central value for $\sigma_{\star}$. Méndez-Abreu et al. (2014), for example, defined any galaxy with $\sigma_{\star}>200\leavevmode\nobreak\ \rm{km}\leavevmode\nobreak\ {s}^{-1}$ to possess a classical bulge. Based on an extensive literature review, Fisher & Drory (2016) updated this value to $\sigma_{\star}>130\leavevmode\nobreak\ \rm{km}\leavevmode\nobreak\ {s}^{-1}$. Other works combine photometric and spectroscopic indicators, for example Neumann et al. (2017) measure the inner slope of the radial velocity dispersion profile within the photometrically-defined bulge region. Several GECKOS galaxies possess a peak in their dispersion profile in their central regions, and for both IC 5244 and IC 1711, this peak value is $>130$ km s-1. Following the literature, we would then attribute these peaks as evidence of classical bulges, however, there are also alternate explanations.
@@
Returning to the cases of IC 1711 and IC 5244, despite their high $\sigma_{\star}$ peaks, we conclude that the BP bulge morphology, clear kinematic nuclear disc signatures, and additional high-order kinematic information point to end-on bars being the likely cause of the $\sigma_{\star}$ behaviour in these galaxies. The observations can be fully explained by the presence of an end-on bar, without the need to invoke significant contribution from a classical bulge. It will be informative to examine the full GECKOS sample for such features, though we note that statistical photometric studies find low fractions of classical bulges at Milky Way stellar masses (e.g. Fisher & Drory, 2011); it may be that the GECKOS sample is simply below the stellar mass range in which we expect to find such features. Indeed, the Milky Way itself possesses cylindrical rotation (Shen et al., 2010) and structural and kinematic properties (Bland-Hawthorn & Gerhard, 2016) inconsistent with the presence of a classical bulge.
@@ 4.3.2 Kinematic disc structures
With this, we have explained the kinematic behaviour of 11/12 GECKOS galaxies (excluding UGC 00903, whose kinematics are considerably more complex and require detailed modelling) by a combination of bars, BP bulges, nuclear discs, and simple disc structure(s). We see no physically-motivated reason to invoke classical bulges in this sample, a conclusion also drawn by Bittner et al. (2020) for the TIMER sample. While certainly not statistical nor representative, the GECKOS sample does span a range of bulge morphologies and bulge-to-total ratios derived from two-component decompositions. The full sample in combination with dynamical modelling will provide excellent insight into the frequency but also necessity (or lack thereof) of compact ($\sim$100s of pc) spheroidal bulges in the local Universe.
@@
ESO 484-036 Yes No Yes Slightly elongated along z=0 line Intermediate
ESO 079-003 Yes No Yes Slightly elongated along z=0 line Intermediate
ESO 120-016 Maybe No Maybe Slightly elongated along z=0 line Intermediate (if barred)
@@
NGC 5775 No No Slightly elongated along z=0 line Unbarred
5 Summary and conclusions
To understand the diversity of kinematic sub-structures in Milky Way-mass galaxies, we examine the stellar kinematics of the first 12 targets observed as part of the GECKOS Survey of nearby edge-on galaxies. After first visually examining optical and mid-IR images to identify BP bulge structures, we examine both the 1D radial profiles and 2D maps of line-of-sight mean velocity ($V_{\star}$), velocity dispersion ($\sigma_{\star}$), and high-order skew ($h_{3}$) and kurtosis ($h_{4}$). Eight of the 12 galaxies possess boxy and/or peanut-shaped central structures, present in either $r$-band optical or 3.6-$\mu$m Spitzer imaging.
@@ 5 Summary and conclusions
  • All galaxies show $h_{3}-V_{\star}$ sign mismatch in outer disc regions consistent with a (quasi-)axisymmetric, rotating disc of stars.
  • All galaxies that show evidence for a clear boxy-peanut shaped structure in images also possess kinematic indicators of a bar. In particular, all of these galaxies show extended regions of $h_{3}$ sign matching with $V_{\star}$ permeating up to 2 kpc off the disc z=0 line. These features are not present in the four non-BP bulge galaxies.
  • Four BP bulge galaxies also host strong kinematic signatures of nuclear discs, including a ‘double hump’ rotation curve, central $\sigma_{\star}$ minima, $h_{3}-V_{\star}$ sign mismatch in the central regions, and strong $h_{4}$ maxima co-spatial with the expected location of the nuclear disc itself.
Comparing to the 2D maps produced by N-body simulations of Iannuzzi & Athanassoula (2015), there is an array of central velocity dispersions, and varying degrees of coherent regions of $h_{3}$-$V_{\star}$ sign match across the bar-dominated regions. In IC 5244, IC 1711, NGC 3957, and PGC 044931, there is strong evidence for the existence of nuclear discs, in the form of central $\sigma_{\star}$ depressions (NGC 3957 and PGC 044931), and central $h_{3}$-$V_{\star}$ sign mismatch (in all four galaxies). Dust obscuration and resolution effects can significantly hinder the recovery of the stellar kinematics in the central regions of edge-on galaxies, so we cannot rule out the possibility that a nuclear disc is present in all of the BP bulge galaxies; we may simply not be able to detect them.
@@ Acknowledgements.
Author contribution statement: This project was devised by AFM, with analysis completed by AFM, and some scripts provided by JvdS. Comments were provided by the whole author list. GECKOS data reduction is led by JvdS and AFM, with assistance from EE, LC, BMC, TM, and BC. nGIST development is led by AFM and JvdS, with contributions and testing from TB, AW, ZW, LSL, LC, FP, MM, DG, EE, and CdSF.
Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 110.24AS. We wish to thank the ESO staff, and in particular the staff at Paranal Observatory, for carrying out the GECKOS observations.
The authors wish to thank the anonymous referee, whose careful reading and interpretation of this manuscript resulted in comments that improved its quality. The authors also wish to thank Adrian Bittner for useful conversations, and his blessing to continue the development of the GIST  pipeline. We also thank Lodovico Coccato for useful conversations relating to MUSE data reduction and analysis, Alex Vazdekis for his permission to distribute the MILES models with the nGIST package, Livia Casagrande whose Honours work provided useful insights, and Michele Cappellari for useful conversations related to pPXF usage.
This research has made use of the NASA/IPAC Extragalactic Database (NED; https://ned.ipac.caltech.edu/) operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research used the Canadian Advanced Network For Astronomy Research (CANFAR) operated in partnership by the Canadian Astronomy Data Centre and The Digital Research Alliance of Canada with support from the National Research Council of Canada the Canadian Space Agency, CANARIE and the Canadian Foundation for Innovation. Part of this research was conducted by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013.
@@
AFM gratefully acknowledges the sponsorship provided by the European Southern Observatory through a research fellowship. DAG is supported by STFC grant ST/X001075/1. TAD acknowledges support from the UK Science and Technology Facilities Council through grants ST/S00033X/1 and ST/W000830/1. FP acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (MCIN/AEI/ 10.13039/501100011033) under grant (PID2021-128131NB-I00) and the European Regional Development Fund (ERDF) ”A way of making Europe”. FP acknowledges support also from the Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant “TraNSLate” No 101108180. FF is supported by a UKRI Future Leaders Fellowship (grant no. MR/X033740/1). AP acknowledges support from the Hintze Family Charitable Foundation. LSL acknowledges CAPES for the support through grant 88887.637633/2021-00 and CNPQ through grant 200469/2022-3 LMV acknowledges support by the German Academic Scholarship Foundation (Studienstiftung des deutschen Volkes) and the Marianne-Plehn-Program of the Elite Network of Bavaria. LC and ABW acknowledge support from the Australian Research Council Discovery Project funding scheme (DP210100337) J.F-B acknowledges support from the PID2022-140869NB-I00 grant from the Spanish Ministry of Science and Innovation. MM acknowledges support from the UK Science and Technology Facilities Council through grant ST/Y002490/1.
@@ Appendix A The nGIST pipeline
The need for further development arose from both the increased computational and scientific requirements of recent IFS surveys. nGIST pushes beyond current limits in terms of memory requirements and deals better with longer optical wavelength ranges and sky residuals that are particularly problematic at redder wavelengths ($\lambda>7000\AA$). Building on the foundations of the GIST pipeline, nGIST also incorporates elements of pipelines developed for recent IFS surveys, including the SAMI stellar kinematics pipeline (van de Sande et al. 2017), and the PHANGS data analysis pipeline (DAP; Emsellem et al. 2022), which in turn is rooted in portions of the MaNGA DAP (Belfiore et al. 2019; Westfall et al. 2019). nGIST represents a significant upgrade in both scientific output and computational performance over GIST.
@@
nGIST commenced from the final release of the GIST  pipeline, v3.1.0888We note that the final release of GIST  requires python $<$ v3.6, which is increasingly deprecated and difficult to install on modern machines via package managers such as conda. At the time of writing, nGIST has been tested up to python v3.11., and has continued to be developed by adding new functionality, and updates that are appropriate for increasingly complex IFS datasets. nGIST is publicly available via a GitHub repository [^9] https://github.com/geckos-survey/ngist, with documentation available [^10] https://geckos-survey.github.io/gist-documentation/. It has been thoroughly tested for MUSE data, but can be employed with any galaxy IFS dataset. Below, we detail the nGIST philosophy, workflow, and major changes introduced since the last version of GIST. For a full description of the original GIST pipeline, we refer the reader to Bittner et al. (2019).
@@
$bias=0.7\times\sqrt{500/N_{\lambda pixels}},$ (3)
@@
In Figure 18, we show the comparison of the stellar kinematic maps from GIST and nGIST for GECKOS iDR1 galaxy IC 1711, with a target $S/N=100$ and minimum spaxel $S/N=5$. For this test, we adopted the v11.0 EMILES Base models with the BaSTI scaled-solar isochrones, that can be used in both GIST and nGIST. For these fits, we removed the ‘unsafe’ SSPs with metallicities [M/H] = -2.27, -1.79, and +0.40. The wavelength range was set to 4800–8900Å, we assumed a constant noise spectrum, and fixed the pPXF bias value to 0.14.
There is a good agreement in Figure 18 for the results from GIST (left column) and nGIST (middle column) for all kinematic parameters, though there are noticeable differences in the outer bins where the contribution from the sky lines and telluric absorption features become more dominant. In particular, $h_{3}$ and $h_{4}$ reach more extreme values in the GIST maps. A one-to-one comparison of the results are presented in Figure 19, where the data are colour-coded by the S/N calculated from the residual of the data minus the best-fit. Here $\sigma_{\star}$ can be seen to exhibit some scatter even at moderate values of S/Nresidual.
@@
nGIST includes an additional spectral cleaning step, and the regions masked as a result of this step are shown as shaded pink regions. While the sky line masks are extensive (based on the UVES sky emission atlas, where every sky line of flux $>5\times 10^{-16}$ erg s-1 Å-1 cm-2 arcsec-2 was masked), some additional masks are always applied by the nGIST cleaning step. The number of masked spectral pixels as a result of the cleaning step increases from central to outer galaxy regions where sky lines begin to dominate. There are clear regions in the outer bin spectrum that benefit from the additional masking, particularly a region around $\lambda\sim 8700\AA$, which the Calcium triplet sits on top of at the redshift of this galaxy. We expect this to be the reason behind the scatter in $\sigma_{\star}$ seen in Figure 19 for a small number of bins. These bins are located around the edges of the galaxy map, where $\sigma_{\star,err}$ is at its highest.
@@
We compare the mean stellar population age and metallicity from GIST and nGIST in Figures 23 and 24. Similar to Pessa et al. (2023), for this comparison we restrict the wavelength region to 4800-7000Å, and adopt a multiplicative polynomial of order 12. To avoid discrepancies in the derived stellar populations resultant from differences in the stellar kinematic solutions, we adopt the Voronoi binning and stellar kinematics solution from the nGIST SPATIAL_BINNING and KIN modules for both. We apply minimal regularisation to the resultant mean stellar age and metallicity maps. The new cleaning algorithm in nGIST has a significant impact on the derived age and metallicity, most noticeably at large radii and scale heights, where the sky line residuals and telluric correction are most pronounced. We also detect a higher metallicity region in the centre of the GIST metallicity map when compared to nGIST  seen in Figure 23. While at first glance, the higher central metallicities of GIST may be more in line with expectations of inside-out disc galaxy growth, we find that the higher metallicity returned is actually due to a poorer fit. Specific faint lines (for example KI at 7664.9Å and 7690.0Å) are particularly prominent in the central regions of IC 1711 (and indeed can be seen in the central spectra at the top of Figure 20), yet not present in the SSP models used. For the GIST pipeline, without automatic masking, these lines bias the best-fit pPXF solution toward higher metallicity, while nGIST’s cleaning algorithm identifies them as significant outliers and masks them, resulting in a lower metallicity estimate. A full test of the impact of clipping outliers and the dependence of wavelength range on the mean stellar population parameters is beyond the scope of this paper, but will be explored in future work.